OpenStudy (goformit100):

Round the number in years and months 18.5

OpenStudy (goformit100):

@Callisto

OpenStudy (goformit100):

Help

OpenStudy (goformit100):

@shadowfiend

OpenStudy (goformit100):

@Mpost1994 plz no don't answer

OpenStudy (anonymous):

19 years and 222 months....don't ask a rhetorical question to be rude and a smart allec please

OpenStudy (goformit100):

@UnkleRhaukus

OpenStudy (goformit100):

Help

OpenStudy (anonymous):

to the nearest what?

OpenStudy (goformit100):

year and month

OpenStudy (anonymous):

well if its 18 years and 5 months then it would be 18 years, but if it were 18.6 then it would be 19 years

OpenStudy (unklerhaukus):

\[18.5[\text{yr}]=18[\text{yr}]+0.5[\text{yr}]\]\[\qquad\qquad =18[\text{yr}]+0.5[\text{yr}]\times\frac{12[\text{mo}]}{[\text{yr}]}\]\[\qquad\qquad =18[\text{yr}]+0.5\times{12[\text{mo}]}\]\[\qquad\qquad=18[\text{yr}]+6{[\text{mo}]}\]

OpenStudy (anonymous):

@UnkleRhaukus so my answer was incorrect, is that right

OpenStudy (goformit100):

@Mpost1994 I WON

OpenStudy (anonymous):

@goformit100 what do you mean you won

OpenStudy (anonymous):

I see conflicting points here

OpenStudy (unklerhaukus):

@heedcom you were trying to round the number 18.5 in years only if you did this correctly you would get 19 years

OpenStudy (goformit100):

0k

OpenStudy (goformit100):

thankyou @UnkleRhaukus

OpenStudy (unklerhaukus):

five always rounds up \(\uparrow\)

OpenStudy (unklerhaukus):

anytime

OpenStudy (anonymous):

but my answer is also correct as I had stated if it were 18 yrs and 5 month its 18 years, so I am also correct @UnkleRhaukus

OpenStudy (anonymous):

I did not say 18.5 is 18years, am I correct @UnkleRhaukus

OpenStudy (unklerhaukus):

oh yeah , in whole numbers \[\frac5{12}= 0.41\dot6=0\]